# AI-Driven Robotics for Free-Space Optics

### v1

@@ Abstract
Tabletop optical experiments are foundational to research in many areas of science, including photonics, quantum optics, materials science, metrology, and biomedical imaging. However these experiments remain fundamentally reliant on manual design, assembly, and alignment—limiting throughput and reproducibility. Optics currently lacks generalizable robotic systems capable of operating across a diverse range of setups in realistic laboratory environments. Here we present OptoMate, an autonomous platform that integrates generative AI, computer vision, and precision robotics to enable automation of free-space optics experiments. Our platform interprets user-defined goals to generate valid optical setups using a fine-tuned large language model (LLM), assembles the setup via robotic pick-and-place with sub-millimeter accuracy, and performs fine alignment using a robot-deployable tool. The system then executes a range of automated measurements—including laser beam characterization, polarization mapping, and spectroscopy tasks. This work demonstrates the first flexible, AI-driven automation platform for optics, offering a path toward remote operation, cloud labs, and high-throughput discovery in the optical sciences.
@@
Figure 1: Overview of OptoMate. A researcher specifies goals for optics experiments, which are interpreted by a large language model (LLM) to generate structured optical setups and executable code. A robotic arm equipped with computer vision then performs optical component identification, pick-and-place operations, positional error correction and fine alignment. Once assembled, the platform performs optical measurements without human intervention.
@@
The precise and repeatable control of experimental workflows is a central challenge across the physical sciences. In recent years, the proliferation of data‑driven methods and large‑scale artificial intelligence (AI) have been transforming how measurements are planned, executed and interpreted [1, 2, 3]. As AI‑enabled approaches become more pervasive, the need for integrated automation in bench‑scale experiments has grown ever more urgent. Pioneering efforts in other fields have already demonstrated the utility of robotics in accelerating scientific progress. In chemistry, automated platforms now execute multi‑step syntheses and characterize reaction outcomes with minimal human intervention, enabling rapid exploration of chemical space [4, 5, 6, 7, 8]. Similarly, programmable fabrication systems have been harnessed to synthesize materials with targeted properties, perform manipulations with atomically precise control and perform characterization tasks [9, 10, 11, 12], while cloud‑based laboratories offer remote access to state-of-the-art instrumentation at large scales [13].
@@
Despite successes in other disciplines, the field of optics has remained predominantly manual in its implementation. While recent studies have explored the use of AI to discover or optimize optical configurations [14, 15, 16, 17], the development of a fully autonomous, end-to-end pipeline from design to implementation remains a challenge. This inertia stems from challenges intrinsic to free-space optical systems, which demand reconfigurability, sub-micron precision in alignment, continuous tuning of sensitive optical elements, and in-situ diagnostics that are highly susceptible to environmental fluctuations. Meeting these stringent demands requires an interdisciplinary approach—one that integrates AI, precision robotics, computer-vision, engineering, and deep expertise in optical systems. A AI-driven robotic framework for optics could overcome these limitations, significantly accelerating experimental workflows. Such a framework would also expand accessibility to complex optical experiments and allow for the exploration of the vast space of optical configurations.
@@
Here, we address this critical gap by introducing OptoMate—an AI-driven robotics platform for free-space optics (Figure 1). OptoMate aims to automate the three fundamental stages of an optical experiment: (1) the design of optical setups, (2) the physical assembly and fine alignment of components, and (3) the execution of measurements. At the design stage, we employ fine-tuned large language models (LLMs) that interpret user-specified goals and generate valid optical setups within the physical constraints of the laboratory environment. These LLM-generated setups are translated into robot control code using a second LLM-based agent. A robotic arm equipped with a computer vision system then performs pick-and-place operations with sub-millimeter precision, followed by fine alignment using a robot-deployable tool. In the final stage, we demonstrate a range of automated measurements commonly performed in optics—including beam characterization, polarization mapping, and transmission spectroscopy.
I Optical setup generation using LLMs
Designing optical setups requires awareness of laboratory spatial constraints, and the ability to translate experimental goals into executable configurations. Traditionally, this process is time-intensive, reliant on expertise, and often involves numerous iterative refinements. As the first stage in the automation pipeline, we present a large language model (LLM)-driven system that automates optical setup design by translating natural language descriptions into valid, spatially-aware, and executable optical setups (Figure 2a).
@@ I Optical setup generation using LLMs
To demonstrate the system’s capabilities, we focus on four canonical optical setups of varying complexity: Michelson interferometer, Mach-Zehnder interferometer, Hong-Ou-Mandel interferometer, and a 4f optical imaging system. Each setup is defined as a list of tuples, with each tuple encoding an optical component ID, x-y position, and orientation. We assume all components to have a fixed z-position (height). From a single reference of each setup, we generate a dataset through geometric data augmentation using similarity transformations—translations, scalings, rotations, and reflections. These synthetic setups are subsequently filtered to ensure physical feasibility: they must lie entirely within the robot-accessible workspace, have no overlapping components, and avoid internal beam paths intersecting the robotic arm’s footprint (for details see Methods).
@@
For each valid setup in the dataset, we generate natural language descriptions using an auto-captioning step. These captions simulate user inputs with varying degrees of specificity, ranging from high-level objectives (e.g., “Give me a Mach-Zehnder interferometer”) to setups with specific requirements (e.g., “Give me a Michelson interferometer setup. The incoming laser beam direction makes an angle of 30 degrees with the x-axis and the setup should have a beamsplitter positioned at approximately (20 cm, 3 cm)”).
@@
We fine-tune a pretrained LLaMA3.1-8B-Instruct model for this task using Quantum-informed Tensor Adaptation (QuanTA) [18], a parameter-efficient fine-tuning (PEFT) method (see Methods). QuanTA utilizes quantum-inspired tensor networks to modify model weights, enabling efficient high-rank adaptations that can capture complex structural relationships while using significantly fewer trainable parameters than full fine-tuning. Our fine-tuning strategy involves two sequential stages: first, we train the model on our augmented dataset to learn the fundamental principles of generating physically valid optical layouts (structure learning); second, we continue the training with prompts containing specific user constraints to enhance the model’s ability to interpret and adhere to precise experimental requirements (user request understanding). Once the LLM generates an optical setup, a validation step evaluates the geometric and physical feasibility of the layout (see Methods). If the configuration fails this validation, it is returned to the model for refinement in a closed-loop correction process. Validated setups are then passed to a secondary LLM-based agent (based on Claude 3.7) which translates the symbolic layout into executable robotic movement commands for physical assembly in the laboratory (see Methods) (Figure 2a).
@@
To assess the effectiveness and necessity of our fine-tuning approach, we compare QuanTA with three alternative strategies: (1) zero-shot inference using GPT-4o, (2) prompt-engineered GPT-4o, and (3) prompt engineering combined with chain-of-thought reasoning using DeepSeek R1 (Figure 2b). We find that our fine-tuned model achieves the highest pre-validation accuracy and the greatest compliance with user-specified constraints. While all approaches ultimately produce a valid setup after iterating through the validation step, the lower pre-validation accuracies of the other methods translate directly into significantly higher computational costs, greater token usage per valid setup, and increased time required to generate a correct solution, compared to fine-tuning.
Figure 2: Optical setup generation using fine-tuned large language models. A researcher provides a natural language input for the desired optical setup. A large language model, fine-tuned using QuanTA [18], generates optical setups. An optical setup includes component IDs, spatial coordinates, and orientations of all components that comprise the setup. This structured output is validated for accuracy and physical feasibility and translated into robot-executable code for assembly. The last panel shows a schematic of a generated setup in the laboratory space, where “R” denotes the footprint of the robotic arm and “WA” indicates the working area of the robotic arm. (b) Comparison of prompting strategies for optical setup generation. Accuracy is reported both with and without user-specified constraints or requests (denoted as “with req.” and “w/o req.”, respectively). Cost and token usage are shown as averages, computed based on the ”with req.” accuracy values (1k = 1,000; \$ = USD). Fine-tuning with QuanTA achieves the highest accuracy with significantly better cost and token efficiency, outperforming chain-of-thought reasoning, prompt engineering, and zero-shot inference.
II Robotic assembly and fine alignment of optical setups
@@
Following the generation of valid and spatially aware optical setups by the LLM-based design system, the next stage involves the autonomous assembly and fine alignment of these setups using a robotic platform. This system is built around a 7-degree-of-freedom robotic arm (UFACTORY xArm7) with measured spatial and angular precisions for placement of $\pm$0.5 mm and 0.2 respectively, a payload capacity of 3.5 kg, and a reach radius of 61 cm. The arm is programmed to perform high-precision placement of optical components, including mirrors, lenses, beam splitters, and detectors. The robotic platform integrates computer vision, LiDAR-based scanning and depth sensing, and custom-designed optical component holders and magnetic bases, all optimized for robotic manipulation.
@@
Each optical component is placed in a 3D-printed housing designed to universally fit standard off-the-shelf optical components and is placed in a 3D-printed base with magnets that provide resistance to unwanted motion during assembly and alignment (Figure 3a). The housing has acArUco fiducial marker affixed to the top surface that facilitates component identification and positioning via stereo 4K visible-light cameras mounted around the workspace. These cameras perform an initial scan of the optical table, allowing the system to estimate component positions and orientations. Once a candidate component is identified, the robotic arm approaches it, and a LIDAR sensor mounted on the end-effector performs a high-resolution scan to refine the object’s pose and geometry. This multi-modal perception pipeline ensures sub-millimeter grasping accuracy and minimizes the risk of collision or misplacement. Upon successful pickup, components are placed at their target positions and orientations according to the layout provided either by a user or by the fine-tuned LLM (see Supplementary Videos 1 and 2). During placement, an error correction step ensure both positional and angular fidelity by rechecking these parameters using the computer vision system.
@@
For optical setups requiring precision beyond the capabilities of initial placement, we develop a robot-deployable fine alignment tool (Figure 3b, also see Supplementary Video 9). This compact, motorized device is constructed to interface with the fine-adjustment knobs located on standard optical mounts, providing electronically controlled, micrometer-scale adjustments to the component’s degrees of freedom. The tool consists of an Arduino Nano microcontroller, a custom PCB with integrated motor drivers, precision-machined Allen keys mounted on low-friction bearing-supported motor shafts, and a dedicated camera and LED illumination system embedded within the housing. For fine alignment tasks, the robotic arm autonomously picks up the fine alignment tool and inserts the Allen keys of the tool into the adjustment knobs of optical components requiring alignment. Accurate insertion is enabled by cameras and LEDs located inside the tool, along with a second ArUco marker on the side surface of the component housing. Alignment adjustments are performed iteratively, guided by feedback from cameras or photodetectors previously placed along the beam path, either through an alignment algorithm or under remote user control. The complete robotic pipeline is shown schematically in Figure 3c (see Methods for more details on the computer vision system, robot pick-and-place operations, custom housing and bases, construction of fine-alignment tool, and correction of positional errors).
@@
To demonstrate the capabilities of the robotic assembly and fine-alignment system, we autonomously construct and align a Michelson interferometer from scratch (Figure 3d, also see Supplementary Videos 3 and 4). The optical layout for the interferometer is generated by the fine-tuned LLM in response to a user-specified constraint on the incoming laser beam direction. Based on the generated design, the computer vision system identifies the required components — two mirrors, a beam splitter, and a camera — and the robotic arm subsequently places them on the optical table (Figure 3d). Following initial assembly, the robotic arm deploys the fine-alignment tool to iteratively adjust the orientation of the mirrors, ensuring the precise overlap of the two beams of the interferometer. Feedback for the alignment is obtained by blocking each arm and centering the beams in the camera’s field of view. The final interferometer and the resulting stable interference pattern are shown in Figure 3d.
@@
Figure 3: Robotic platform for autonomous assembly and alignment of optical setups. (a) Schematic of the robotic platform, consisting of a 7-degree-of-freedom robotic arm equipped with a gripper and a LiDAR camera, as well as external 4K cameras mounted around the workspace to provide stereo vision. Optical components are housed in custom 3D-printed casings labeled with ArUco fiducial markers for identification and positioning. The components are placed in 3D-printed magnetic bases that resist unwanted motion during handling. (b) A schematic of the robot-deployable fine-alignment tool, which consists of two motors with Allen keys mounted on their shafts and an onboard computer vision system enabled by a camera and a white-light LED. (c) Pipeline for autonomous assembly of optical setups. First, the computer vision system performs a coarse estimation of the location, orientation, and angles of the components in the working area using the 4K cameras. The robotic arm then refines this estimate using the onboard LiDAR sensor. Next, a coarse pick-and-place operation positions and orients the component as desired within the optical setup, achieving sub-millimeter and sub-degree precision. Finally, the robot-deployable fine-alignment tool executes micro-adjustments, achieving the sub-arcminute angular precision required for functional optical alignment. This multi-stage pipeline enables robust and repeatable assembly of complex free-space optical configurations. (d) Autonomously assembled and aligned Michelson interferometer setup in the laboratory. The insets show a top-down view of the interferometer and a camera image of the achieved interference pattern.

### v2

# AI-Driven Robotics for Optics

@@ Abstract
Optics is foundational to research in many areas of science and engineering, including nanophotonics, quantum information, materials science, biomedical imaging, and metrology. However, the design, assembly, and alignment of optical experiments remain predominantly manual, limiting throughput and reproducibility. Automating such experiments is challenging due to the strict, non-negotiable precision requirements and the diversity of optical configurations found in typical laboratories. Here, we introduce a platform that integrates generative artificial intelligence, computer vision, and robotics to automate free-space optical experiments. The platform translates user-defined goals into valid optical configurations, assembles them using a robotic arm, and performs micrometer-scale fine alignment using a robot-deployable tool. It then executes a range of automated measurements, including beam characterization, polarization mapping, and spectroscopy, with consistency surpassing that of human operators. This work demonstrates the first flexible, AI-driven automation platform for optics, offering a path towards remote operation, cloud labs, and high-throughput discovery in the optical sciences.
@@
Figure 1: Overview of the AI-driven robotics platform for free-space optics. A researcher specifies goals for optics experiments, which are interpreted by a large language model (LLM) fine tuned using QuanTA [1] to generate an optical setup design (“Optics agent”). The setup is then validated for accuracy and physical feasibility and translated into robot-executable code for assembly (“Coding agent”). Subsequently, a robotic arm equipped with computer vision performs optical component identification, pick-and-place operations, positional error correction, and fine alignment. Once assembled, the platform autonomously performs optical measurements such as characterization and imaging tasks.
@@
Precise and repeatable control of experimental workflows is a central challenge across the physical sciences. In recent years, the proliferation of data‑driven methods and large‑scale artificial intelligence (AI) has been transforming how measurements are planned, executed, and interpreted [2, 3, 4]. As AI‑enabled approaches become more pervasive, the need for integrated automation in bench‑scale experiments has grown ever more urgent. Pioneering efforts in many fields have already demonstrated the utility of robotics in accelerating scientific progress. In chemistry, automated platforms now execute multi‑step syntheses and characterize reaction outcomes with minimal human intervention, enabling rapid exploration of chemical space [5, 6, 7, 8, 9, 10, 11]. Similarly, programmable fabrication systems have been harnessed to synthesize materials with targeted properties, perform manipulations with atomically precise control, and perform characterization tasks [12, 13, 14, 15, 16], while cloud‑based laboratories offer remote access to state-of-the-art instrumentation at large scales [17].
@@
Optics is both a fundamental science, concerned with the study of light and its interactions with matter, and an enabling technology that permeates nearly every scientific domain. It provides the basis for instrumentation in observational astronomy, advanced microscopy techniques in biology, spectroscopic and imaging tools in materials research, and precision experiments in atomic, molecular, and chemical physics. Despite successes in other disciplines, the field of optics has remained predominantly manual in its implementation. While recent studies have explored the use of AI to discover or optimize optical configurations [18, 19, 20, 21], the development of a fully autonomous, end-to-end pipeline from design to implementation remains an open problem. This inertia stems from major challenges intrinsic to free-space optical systems, which demand reconfigurability, sub-micron precision in alignment, continuous tuning of sensitive optical elements, and in-situ diagnostics that are highly susceptible to environmental fluctuations. Meeting these stringent demands requires an interdisciplinary approach: one that integrates AI, precision robotics, computer vision, engineering, and deep expertise in optical systems. An AI-driven robotic framework for optics automation could overcome these limitations, significantly accelerating experimental workflows. Such a framework would also expand accessibility to complex optical experiments and allow for the exploration of the vast space of optical configurations directly in the laboratory.
@@
Here, we address this critical gap by introducing the first AI-driven robotics platform for free-space optics (Figure 1) where we aim to automate the three fundamental stages of an optical experiment: (1) design, (2) physical assembly and fine alignment, and (3) execution of measurements. At the design stage, we employ fine-tuned large language models (LLMs) that interpret user-specified goals and generate valid optical setups within the physical constraints of the laboratory environment. These LLM-generated setups are then translated into robot control code using a second LLM-based agent. A robotic arm equipped with a computer vision system performs pick-and-place operations with sub-millimeter precision, followed by fine alignment using a purpose-built robot-deployable tool. In the final stage, we demonstrate a range of automated measurements commonly performed in optics laboratories, including beam characterization, polarization mapping, transmission spectroscopy, and real-space setup optimization.
I Optical setup design using LLMs
Designing free-space optical experiments requires translating high-level experimental intent into concrete spatial configurations of optical components. This requires reasoning about laboratory constraints and propagation paths. Humans accomplish this by combining domain knowledge of optical physics with spatial intuition—a capability that has remained fundamentally inaccessible to automated systems. To bridge this gap, we adopt a generative approach based on large language models (LLMs) fine-tuned to generate optical setups, enabling them to produce physically valid and robot-executable experimental designs from natural language descriptions (Figure 1).
@@ I Optical setup design using LLMs
To demonstrate the system’s capabilities, we focus on four canonical optical setups of varying complexity: Michelson interferometer, Mach-Zehnder interferometer, Hong-Ou-Mandel interferometer, and a 4f optical imaging system. We define an optical setup as a list of tuples, with each tuple encoding an optical component identifier, x-y position, and orientation. For simplicity, we assume all components to have a fixed z-position (height). From a single reference of each setup, we generate a dataset through geometric data augmentation using similarity transformations (translations, scaling, rotations, and reflections). These synthetic setups are subsequently filtered to ensure physical feasibility: they must lie entirely within the robot-accessible workspace, have no overlapping components, and avoid internal beam paths intersecting the robotic arm’s footprint (see Methods for details).
@@
For each setup in the dataset, we generate natural language descriptions using an auto-captioning step. These captions simulate user inputs with varying degrees of specificity, ranging from high-level objectives (e.g., “Give me a Mach-Zehnder interferometer”) to setups with specific requests (e.g., “Give me a Michelson interferometer setup. The incoming laser beam direction makes an angle of 30 degrees with the x-axis and the setup should have a beam splitter positioned at approximately (20 cm, 3 cm)”). Together, these steps produce a scalable labeled dataset of physically valid, linguistically varied examples from which the model can learn to generalize, without the need for expensive optical simulation.
@@
We then fine-tune a pretrained LLaMA3.1-8B-Instruct model for this task using Quantum-informed Tensor Adaptation (QuanTA) [1], a parameter-efficient fine-tuning (PEFT) method (see Methods). QuanTA utilizes quantum-inspired tensor networks to modify model weights, enabling efficient high-rank adaptations that can capture complex structural relationships while using significantly fewer trainable parameters than full fine-tuning. Our fine-tuning strategy involves two sequential stages: first, we train the model on our augmented dataset to learn the fundamental principles of generating physically valid optical layouts (structure learning); second, we continue the training with prompts containing specific user constraints/requests to enhance the model’s ability to interpret and adhere to precise experimental requirements (user request understanding). Once the LLM generates an optical setup, a validation step evaluates the geometric and physical feasibility of the layout (see Methods). If the configuration fails this validation, it is returned to the model for refinement in a closed-loop correction process. Validated setups are then passed to a secondary LLM-based agent (based on Claude 3.7) which translates the symbolic layout into executable robotic movement commands for physical assembly in the laboratory (see Methods).
@@
To assess the performance of our fine-tuning approach, we compare QuanTA with three alternative strategies: (1) zero-shot inference using GPT-4o, (2) prompt-engineered GPT-4o, and (3) prompt engineering combined with chain-of-thought reasoning using DeepSeek-R1 (Figure 2). We find that our fine-tuned model achieves the highest pre-validation accuracy ($80\%$), defined as the fraction of correct setups over total attempts. Under user-specified constraints, such as exact placement of some components, the model maintains high accuracy while also achieving strong compliance (placement within 1 mm and 0.1 of the requested position), whereas other strategies, though compliant, fail to yield valid setups (Figure 2a,b). While all approaches would eventually converge to a valid setup after enough iterations through the validation step, the significantly higher accuracy of our fine-tuned model directly translates into reduced computational costs, lower token usage per valid setup, and shorter times to obtain a correct solution (Figure 2c).
Figure 2: Comparison of prompting strategies for autonomous optical setup generation. (a) Accuracy across prompting strategies, averaged over all four setup types. (b) User compliance rates, reported as the fraction of setups meeting the user-requested conditions, regardless of the validity of setups. (c) Average token usage (log scale) per valid setup generated for each prompting strategy.
II Robotic assembly and fine alignment
@@
Following the generation of valid and spatially aware optical setups by the LLM-based design system, the next stage involves the autonomous assembly and fine alignment of these setups using a robotic platform. This system is built around a 7-degree-of-freedom robotic arm (UFACTORY xArm7, see Methods) and is deployed in a standard optics laboratory. The robotic arm is programmed to perform high-precision placement of optical components, including mirrors, lenses, beam splitters, and detectors. We integrate the robotic arm with computer vision, LiDAR-based scanning and depth sensing, and custom-designed optical component holders and magnetic bases, all optimized for robotic manipulation.
Each optical component is placed in a 3D-printed housing that is designed to universally fit standard off-the-shelf optical components. The component is secured to a 3D-printed magnetic base that provides resistance to unwanted motion during assembly and alignment (Figure 3a). The housing has an ArUco fiducial marker affixed to the top surface that allows for component identification and positioning via stereo 4K visible-light cameras mounted around the workspace. These cameras perform an initial scan of the optical table, allowing the system to estimate component positions and orientations.
@@
Once a candidate component is identified, the robotic arm approaches it, and a LIDAR sensor mounted on the end-effector performs a high-resolution scan to refine the object’s pose and geometry. This multi-modal perception pipeline ensures sub-millimeter grasping accuracy and minimizes the risk of collision or misplacement. Upon successful pickup, components are placed at their target positions and orientations according to the layout provided either by a user or by the fine-tuned LLM (see Supplementary Videos 1 and 2). During placement, an error correction step ensures both positional and angular fidelity by rechecking these parameters using the computer vision system.
@@
For optical setups requiring precision beyond the capabilities of the initial placement, we developed a robot-deployable fine alignment tool (Figure 3b, also see Supplementary Video 9). This compact, motorized device is constructed to interface with the fine-adjustment knobs located on standard optical mounts, providing electronically controlled, micrometer-scale adjustments to the component’s degrees of freedom. The tool consists of an Arduino Nano microcontroller, a custom PCB with integrated motor drivers, Allen keys mounted on motor shafts, and a dedicated camera and LED illumination system. For fine alignment tasks, the robotic arm autonomously picks up the fine alignment tool and inserts the Allen keys of the tool into the adjustment knobs of optical components. Alignment adjustments are performed iteratively, guided by feedback from cameras or photodetectors previously placed along the beam path. The complete robotic pipeline is shown schematically in Figure 3c (see Methods for more details on the computer vision system, robot pick-and-place operations, custom housing and bases, construction of fine-alignment tool, and correction of positional errors).
@@
To demonstrate the capabilities of the robotic assembly and fine-alignment system, we first ask the AI-driven robotic system to autonomously assemble and align a Michelson interferometer from scratch (Figure 3d, Supplementary Videos 3 and 4). The optical layout for the interferometer is generated by the fine-tuned LLM in response to a user-specified constraint on the incoming laser beam direction. Based on the generated design, the computer vision system identifies the required components — two mirrors, a beam splitter, and a camera — and the robotic arm subsequently places them on the optical table (Figure 3d). Following initial assembly, the robotic arm deploys the fine-alignment tool to adjust the orientation of the mirrors, ensuring the precise overlap of the two beams of the interferometer. Feedback for the alignment is obtained by blocking each arm and centering the beams in the camera’s field of view. The final interferometer and the resulting stable interference pattern are shown in Figure 3d.
As a second demonstration, we ask the AI-driven robotic system to autonomously assemble a 4f beam cleaning setup consisting of two lenses and a pinhole. As before, the optical layout is generated by the fine-tuned LLM after specifying the incoming laser direction. The robotic arm places the components to create a 4f setup, and a camera images the beam before and after setup assembly (Figure 3e, also see Supplementary Video 10).
@@
Figure 3: Autonomous assembly and alignment of optical setups. (a) Schematic of the robotic platform, consisting of a 7-degree-of-freedom robotic arm equipped with a gripper and a LiDAR camera, as well as external 4K cameras mounted around the workspace to provide stereo vision. Optical components are housed in custom 3D-printed casings labeled with ArUco fiducial markers for identification and positioning. The components are placed in 3D-printed magnetic bases that resist unwanted motion during handling. (b) Schematic of the robot-deployable fine-alignment tool, which consists of two motors with Allen keys mounted on their shafts and an onboard computer vision system enabled by a camera and a white-light LED. (c) Pipeline for autonomous assembly of optical setups. First, the computer vision system performs a coarse estimation of the position and orientation of the components in the working area using the 4K cameras. The robotic arm then refines this estimate using the onboard LiDAR sensor. Next, a coarse pick-and-place operation positions and orients the component as desired within the optical setup, achieving sub-millimeter and sub-degree precision. Finally, the robot-deployable fine-alignment tool executes micro-adjustments, achieving the sub-arcminute angular precision required for functional optical alignment. This multi-stage pipeline enables robust and repeatable assembly of complex free-space optical configurations. (d) Autonomously assembled and aligned Michelson interferometer setup in the laboratory. The insets show a top-down view of the interferometer and a camera image of the achieved interference pattern. (e) Autonomously assembled beam cleaning setup (4f optical system) in the laboratory. The insets show the beam profile before and after passing through the setup.